Racing greyhounds are capable of rounding corners at very high speeds. A typical greyhound track has turns that are 45 m diameter semicircles. A greyhound can run around these turns at a constant speed of 14 m/s .
Answer #1

acceeleration = v2/R

a = 142/(45/2) = 8.711 m/s^2

a = 8.711 / 9.81 = 0.89 g

So acceleration of the car is 8.711 m/s^2